6 7 Question 1 (30 points): Consider an escalation rate of 11% for all costs (including environmental remediation cost) and 8If using excel please show equations used/explanation
Answer #1

Esc. rate 2 8% Inflow Outflow 3 11% 5 Year 1 10 11 12 cash inflow cash out flow -54000 13000 -16000 12000 10000 10000 9000 90

Formulas used :-

Escalated cash inflow =PV($C$2,A8,0,-B8) (and drag it to the other cells)

Escalated cash out flow=PV($C$3,A7,0,-C7)

Net Cashflow =B7+C7​​​​​C7

Escalated IRR =IRR(F7:F18)

Escalated NPV=NPV(12%,F7:F18)

11th Year cashi flows includes 6000 salvage value and cash out flow is remediation cost.

I hope my efforts will be fruitful to you...?